OpenStudy (anonymous):

what is the pH of 100 mM HCl?

OpenStudy (abb0t):

100mM = 10*10^-4 OR 1*10^-3 [H+] = 1*10^-3 pH = -log 1*10^-3 pH = 6.90

OpenStudy (cwrw238):

? - how did you get 6.9???

OpenStudy (cwrw238):

shouldnt it be 3?

OpenStudy (abb0t):

Oh, you're right. It should be 3. My mistake. I should of known it was wrong. a pH of almost 7 for HCl makes no sense :P

OpenStudy (anonymous):

I don't get why it should be three - surely the pH is the -log10 of the H+ concentration so it's log10(0.1) =-1 so pH = 1. If it was 1M HCl the pH woould be 0, since log 10 of 1 is 0. pH=3 if the concentration of H+ is 1mM surely?